2. (a) *D1* Let A be a fluctuating quantity like the position, velocity, or energy of a molecule. At any given time, we can quantify the fluctuation of A away from the average (A) by the quantity δ = A-(A). Show that the mean square fluctuation in A can be written in terms of its mean and mean square values. «6A)2-(A2>-(A)2. The following properties of averages might be useful (X and Y are fluctuating quantities, c is a constant) · 〈XY〉 = 〈X) 〈Y) only if X and Y are not correlated. .e., the outcome of X does not depend on Yand vice versa (2 pts) (b) *D1* Generalize this result to the case of two different fluctuating quantities A and B, i.e., show that 〈6A8B〉 = 〈AB) _ 〈A)(B). (2 pts) *D1* quantities A and B. To see this, consider tossing a pair of coins A and B. We will assign values of-1 and 1 for heads and tails, respectively. The coins are fair, so 〈A)-〈B) 0. Calculate (6ASB) for the following cases: (c) The quantity 46B) is often used to measure the correlation between two fluctuating i. A and B are statistically independent, i.e., they behave like normal coins ii. By some dark magic, B always shows the same result as A iii. By some even darker magic, B always shows a result opposite to A Explain in what sense the values of 〈5.45B) can be interpreted as a measure of correlation. (2 ts
Answer #1

left langle delta A^2 ight angle=langle(A-langle A angle)^2 angle=langle A^2+langle A angle^2-2Alangle A angle angle

langle A angle acts like a constant, then

left langle delta A^2 ight angle=langle A^2 angle+langle A angle^2-2langle A anglelangle A angle=langle A^2 angle-langle A angle^2

left langle delta Adelta B ight angle=langle (A-langle A angle)( B-langle B angle) angle

left langle delta Adelta B ight angle=langle AB-langle A angle B-Alangle B angle+langle A anglelangle B angle angle

Again langle A angle,langle B angleacts like a constant, then

left langle delta Adelta B ight angle=langle AB angle-langle A anglelangle B angle

(c)

(i) When A and B are independent,

left langle delta Adelta B ight angle=langle AB angle-langle A anglelangle B angle=langle A anglelangle B angle-langle A anglelangle B angle=0

(ii)

(AB) × 1 × 1+ ×(-1) × (-1) = 1 2

Thus,

langle delta Adelta B angle=1

(iii)

(AB) =-× 1 × (-1) +ー× (-1) × 1 =-1

Thus,

langle delta Adelta B angle=-1

The cases (i),(ii) and (iii) correspond to uncorrelated, 100% correlated and 100% anti-correlated events A and B.